OpenStudy (anonymous):

the price of my DSL service went from $40 per moth to $45 per month. This reflects what percent increase in monthly cost?

OpenStudy (mathwizz):

i think it would be 5% becaus percent means out of 100 and adding five is like adding 5%

OpenStudy (mathwizz):

do u understand

OpenStudy (anonymous):

yes i think so

OpenStudy (mathwizz):

ok. great

OpenStudy (anonymous):

so would the equation be x/100=?????

OpenStudy (mathwizz):

it would be c=40+5n.

OpenStudy (anonymous):

o okay thanks!

OpenStudy (mathwizz):

no it would be c=40+45n